OpenStudy (anonymous):

evaluate limit lim as x approaches 2 of (x-2)/(Sqrt x)-(Sqrt (4-x))

OpenStudy (anonymous):

\[(x-2)\over \sqrt{x}-\sqrt{4-x}\] right?

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

o sorry

OpenStudy (anonymous):

Did I rewrite the problem correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

When you plug in 2 you get \[0 \over 0\], a indeterminate. SO we may apply L'Hopital Rule

OpenStudy (anonymous):

or rationalize the denominator if you have not gotten to l'hopital yet

OpenStudy (anonymous):

then cancel, plug in 2, and get the answer

OpenStudy (anonymous):

i will write it if you like

OpenStudy (anonymous):

manny are familiar with LHopital rule?

OpenStudy (anonymous):

\[\frac{x-2}{\sqrt{x}-\sqrt{4-x}}=\frac{x-2}{\sqrt{x}-\sqrt{4-x}}\times \frac{\sqrt{x}+\sqrt{4-x}}{\sqrt{x}+\sqrt{4-x}}\]

OpenStudy (anonymous):

sqrt(2) ans...

OpenStudy (anonymous):

sqrt(2)= a.41421

OpenStudy (anonymous):

\[=\frac{(x-2)(\sqrt{x}+\sqrt{x-4})}{2x-4}\]

OpenStudy (anonymous):

\[=\frac{\sqrt{x}+\sqrt{4-x}}{2}\]

OpenStudy (anonymous):

now replace x by 2 and get \[\frac{2\sqrt{2}}{2}=\sqrt{2}\]

OpenStudy (anonymous):

ty for the answers